Questions on APPLICATIONS OF DERIVATIVES.......
Find the equation of the normal to the curve 
y= (1+x)y + sin-1(sin2x) at x=0???
ITS IITJEE(1993) QUESTION...
CAN SOMEBODY PLEASE HELP ME???

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
at x = 0 in this curve we get
y = 1^y
 
hence y = 1
 
now we have to draw a normal at (0, 1)
 
now dy/dx at (0, 1) = 1
 
hence slope of normal = – 1
 
now equation of normal = 
 
y – 1 = – 1( x – 0)
 
y – 1 = – x
x + y = 1
 
Regards
Arun (askIITians forum expert)